Present instrument status
Since its re-deployment at El Leoncito
in February 2006, the Argentine airglow spectrometer
has completed six more years of remarkably successful data acquisition. The
instrument presently looks like this
(status: 2 Dec 2011).
In early 2003, several technical problems had piled up that could not be
repaired on site and led to abortion of data acquisition. The instrument was
transported to Buenos Aires and has undergone a number of modifications,
upgrades, and refurbishments in our laboratory at IAFE.
September 15, 2009, we can now access the latest data each day, via the
internet. However, processing and uploading to WDC/DLR is normally done on a weekly schedule.
The following faults were diagnosed:
- Elastic adhesive coupling in micrometer head drive broken
- Filter tilt unit axis block glue broken
- Binary port input from upper tilt limit sensor dead
- Freon escapes from solder leaks in evaporator part of detector cooling
- Supply line to last PM dynode open contact (only capacitive voltage
The following diagnostics, repairs and upgrades were performed:
- Mechanical problem of filter tilt mechanism was caused by electronic
- Input port failure has been repaired.
- The two broken couplings in tilting filter drive were repaired.
- Inspection of the micrometer screw and contact steel ball showed
very little wear.
- Traces of wear on the micrometer head tip were retouched / steel ball
- DIN connectors were replaced by rugged MIL STD
versions; cabling layout improved.
- Filter tilt, pulse amplification, and counting systems again fully
- 16 MHz 286/7 motherboard replaced by 20 MHz 386/7.
- Replacement of hard disk after partial failure, with complete recovery
of contents and directory structure.
- Failed input port on controller card repaired by installing additional
74LS244 IC (with some spare ports).
- Installation of stepping motor control unit
directly in distribution box, eliminating several cables/connectors.
- Detection and repair of freon leak in evaporator (union of entrance
capillary and evap coil).
- Detection and repair of second (principal) freon leak in capillary;
resoldered with copper sleeve.
- Successful cooling test
- Long bolts put on bottom plate as stands for handling
- Mechanical support added to high voltage supply resistor in preamp
- Replacement photomultiplier installed, optical alignment checked.
- Photomultiplier dark count checked and found very low.
- Neon spectra taken to check spectral scan range.
- Instrument spectral calibration
- New wavelength calibration included in control software.
- Last dynode contact problem repaired.
- Instrument transported to site
- El Leoncito instrument setup and
performance check campaign, Feb 3 - 7,
- A simple daylight sensor (sky-looking photoresistor connected to the PC
printer port) avoids daytime measurement in case of instrument clock
failure (also for monitoring RTC drift, weather conditions, etc.)
- The internet server for data access (since September 2009)
is now used for RTC time correction.
Last update: Jan 13, 2012
IAFE Aeronomy Group